This is a problem I am having a hard time doing. log(x+9)=1-logx
Answer by RAY100(1637)

log(x+9)=1-logx
log(x+9) +logx=1
to multiply is to add logs
log(x+9)*(x)=1
log(x^2+9x)=1
raise both sides as power of 10
10^log(x^2+9x)=10^1
(x^2+9x) =10
x^2+9x-10=0
(x+10)(x-1)=0
x=1, and -10 (not realistic answer, no log(- number)
check x=1
log(1+9)=1-log1
log10=1-log1
1=1-0 ok